# Sample Complexity of Diffusion Model Training Without Empirical Risk Minimizer Access

### v2

@@ Abstract
Diffusion models have demonstrated state-of-the-art performance across vision, language, and scientific domains. Despite their empirical success, prior theoretical analyses of the sample complexity suffer from poor scaling with input data dimension or rely on unrealistic assumptions such as access to exact empirical risk minimizers. In this work, we provide a principled analysis of score estimation, establishing a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$. Our approach leverages a structured decomposition of the score estimation error into statistical, approximation, and optimization errors, enabling us to eliminate the exponential dependence on neural network parameters that arises in prior analyses. It is the first such result which achieves sample complexity bounds without assuming access to the empirical risk minimizer of score function estimation loss.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks [38, 4], text generation [25], sequential data modeling [37], reinforcement learning and control [42], and life sciences [19, 30]. For a more comprehensive exposition of applications, we refer readers to survey paper [9].
@@
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. The iteration complexity or convergence has been studied in [23, 5, 24, 18, 11, 26, 27], while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function. We note that the [40, 39, 32, 8] have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, [15] did an analysis as a baseline for the approach in [7], and showed that this approach can have polynomial dependence in $\epsilon$, but have exponential dependence on neural network parameters, making the result less important for large neural networks. The key previous work without an exponential dependence on neural network or dimension is [15] which obtains $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity [^1] We note that [15] claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix B. . However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss a significant restriction that was explicitly highlighted as an open problem in [15]. In this paper, we address this open problem and overcome this limitation by establishing a sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-6})$ without assuming access to the ERM of the score estimation loss. This represents a key step toward bridging the gap between the theory and practice of diffusion models. Specifically, we address the following fundamental question:
@@
How many samples are required for a sufficiently expressive neural network to estimate the score function well enough to generate high-quality samples using a DDPM-type algorithm, without assuming access to the empirical minimizer of the score function estimation loss?
In this paper, we provide a refined analysis that yields a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$, while avoiding the exponential dependence on data dimension and neural network parameters as in [15]. This is achieved through a decomposition of the total variation error into distinct terms associated with forward process approximation, discretization of the reverse-time process, and score estimation. Using our novel analysis of the score estimation error, we obtain the sample complexity bounds without exponential dependence on data dimension. Our principled analysis accounts for both statistical and optimization errors while not assuming the access to the empirical risk minimizer of the score estimation loss, and achieving above mentioned sample complexity bounds.
@@
Exponential Dependence
on Data Dimension
[40] $\tilde{O}\left(\epsilon^{-d}\right)$ Yes Yes
[39] $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes Yes
[32] $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes Yes
[8] $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes Yes
[15]1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes No
This work $\tilde{O}(\epsilon^{-6})$ No No
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. The additional $\tilde{\mathcal{O}}(\epsilon^{-1})$ factor in our sample complexity bound arises from relaxing the ERM assumption on the score estimation loss. This should not be interpreted as a weaker result, but rather as a reflection of the inherent trade-off between making the ERM assumption and the resulting sample complexity bounds. Here $d$ represents the data dimension. For further details on how the sample complexity bounds are derived for [15], see Appendix B.
The score estimation is bounded to account for approximation error due to the limited expressiveness of the function class, statistical error due to finite sample size, and the optimization error due to not reaching global minimum during training. We note that the dimension-dependent term is the approximation error, which can be bounded for over-parametrized neural networks. This work provides the first analysis in this area that explicitly quantifies the optimization error, introducing a novel recursive analysis of the error at each stochastic gradient descent (SGD) step. This approach captures the effect of using a finite number of SGD iterations to estimate the score function.
Furthermore, this work introduces a novel analysis of the statistical error that avoids the exponential dependence on neural network parameters observed in [7], and does not rely on the quantile-based formulation used in [15]. This is achieved by directly bounding the gap between the expected and empirical loss over a finite function class, thereby sidestepping the need for bounds that scale poorly with network size. A key step in our analysis leverages the quadratic growth property implied by the Polyak-Łojasiewicz (PL) condition assumed in Assumption 2, which governs the score estimation error. This condition plays a critical role in controlling the sample complexity. While stronger growth conditions (e.g., lower Lipshitz) could in principle yield a sample complexity of $\widetilde{\mathcal{O}}(\epsilon^{-3})$, standard neural networks typically satisfy only the PL condition. Moreover, recent work [15] assumes access to the empirical risk minimizer (ERM) of the score estimation loss an assumption that is unrealistic in practice. Under this assumption, they achieve a sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-5})$. In contrast, by removing this impractical assumption and instead relying on the Polyak-Łojasiewicz (PL) condition, our analysis yields a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$. This highlights a trade-off between assuming access to an ERM and the resulting sample complexity: relaxing the ERM assumption comes at the cost of a slightly worse (yet more realistic) sample bound.
@@
  • Finite time sample complexity bounds. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-6})$ for score-based diffusion models, and do not have exponential dependence on the data dimension or neural network parameters. Our analysis avoids the unrealistic assumptions used in prior work such as access to empirical loss minimizer.
@@
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in [10] showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works [23, 5, 24, 18, 11, 26, 27]. In particular, [5, 23] establish iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates [22, 27, 11]. Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in [24, 18, 26]. In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the error introduced by neural network approximation, data sampling, and optimization are also accounted.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in [8, 40, 39, 32]. However, these bounds are exponential in the data dimension. Further, authors of [15] use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
@@ 2 Preliminaries and Problem Formulation
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process [33], defined by the following SDE
@@
where $B_{t}$ denotes standard Brownian motion. The solution of this SDE in the closed form is given by
$x_{t}\sim e^{-t}x_{0}+\mathcal{N}(0,\sigma_{t}^{2}I),\quad\text{with }\sigma_{t}^{2}=1-e^{-2t}.$ (2)
@@
This is typically achieved using stochastic time-reversal theory [3], which yields a corresponding reverse-time SDE as follows.
@@
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions.
Assuming the second moment of $p_{0}$ is bounded, i.e., $m_{2}^{2}:=\mathbb{E}_{x\sim p_{0}}[\|x\|^{2}]$, it is sufficient to choose $T=\mathcal{O}(\log(m_{2}^{2}/\epsilon))$ to ensure that the reverse process generates samples within $\epsilon$-total variation distance of $p_{0}$ [6].
In practice, the continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework [16], where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix C. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
Problem formulation: Let the score function be approximated using a parameterized family of functions $\mathcal{F}_{\Theta}=\{s_{\theta}:\theta\in\Theta\}$, where each $s_{\theta}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ is typically represented by a neural network of depth $D$ and width $W$ with smooth activation functions. Given $n$ i.i.d. samples $\{x_{i}\}_{i=1}^{n}$ from the data distribution $p_{t}$, the score network is trained by minimizing the following time-indexed loss:
$\mathcal{L}_{t}(\theta):=\mathbb{E}_{x_{t}\sim p_{t}}\left[\|s_{\theta}(x_{t},t)-\nabla\log p_{t}(x_{t})\|^{2}\right].$ (4)
Let $p$ be the original data distribution. We additionally define the following probability distributions:
$p_{t_{0}}\coloneqq$ Distribution obtained after backward diffusion till time $t_{0}$ steps starting form $p_{T}$
$p^{dis}_{t_{0}}\coloneqq$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
@@
Objective. Our goal is to quantify how well the learned generative model $\hat{p}_{t_{0}}$ approximates the true data distribution $p$ in terms of total variation (TV) distance. Specifically, we aim to show the number of samples needed so that with high probability, the TV distance $\mathrm{TV}(p,\hat{p}_{t_{0}})$ is bounded by $\mathcal{O}(\epsilon)$, where $\epsilon$ is the $L^{2}$ estimation error of the score function.
@@ 3 Sample Complexity of Diffusion Models
We first state the assumptions required throughout this work. We begin with a commonly used smoothness condition [17, 21] on the score function, which ensures well-behaved gradients and facilitates convergence guarantees for stochastic first-order methods.
@@
For all $t\in[0,T]$, the population loss $\mathcal{L}_{t}(\theta)$ is $\kappa$-smooth with respect to the parameters $\theta$, i.e., for all $\theta,\theta^{\prime}\in\Theta$
$\|\nabla\mathcal{L}_{t}(\theta)-\nabla\mathcal{L}_{t}(\theta^{\prime})\|\leq\kappa\|\theta-\theta^{\prime}\|.$ (5)
We assume that the estimators of the gradients ${\nabla}\mathcal{L}_{t}(\theta)$ have bounded variance.
@@
The bounded gradient variance assumption is a standard assumption used in SGD convergence results such as [20] and [1]. This assumption holds for a broad class of neural networks under mild conditions, such as Lipschitz-continuous activations (e.g., GELU), bounded inputs, and standard initialization schemes [2, 12]. The smoothness condition ensures that the gradients of the loss function do not change abruptly, which is crucial for stable updates during optimization. This stability facilitates a controlled optimization dynamics when using first-order methods like SGD or Adam, ensuring that the learning process proceeds in a well-behaved manner. We emphasize that this smoothness assumption and bounded gradient assumptions are not present in prior works such as [15] and [7], which evaluate the sample complexity of diffusion models under the assumption of access to the empirical risk minimizer, rather than SGD based optimization.
Assumption 2 (Polyak - Łojasiewicz (PL) condition.).
The loss $\mathcal{L}_{t}(\theta)$ for all $t\in[0,T]$ satisfies the Polyak–Łojasiewicz condition, i.e., there exists a constant $\mu_{t}>0$ such that
$\frac{1}{2}\|\nabla\mathcal{L}_{t}(\theta)\|^{2}\geq\mu_{t}\left(\mathcal{L}_{t}(\theta)-\mathcal{L}_{t}(\theta^{*})\right),\quad\forall~{}\theta\in\Theta,$ (7)
where $\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}_{t}(\theta)$ denotes the global minimizer of the population loss.
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses [29]. Prior works such as [15] and [7] implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in [15] and the definition of $\hat{f}$ in Theorem 13 of [7]). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.

### v5

# Improved Sample Complexity For Diffusion Model Training Without Empirical Risk Minimizer Access

@@ Abstract
Diffusion models have demonstrated remarkable performance in generating high-dimensional samples across domains such as vision, language, and the sciences. Although continuous-state diffusion models have been extensively studied both empirically and theoretically, discrete-state diffusion models, essential for applications involving text, sequences, and combinatorial structures, they remain significantly less understood from a theoretical standpoint. In particular, all existing analyses of discrete-state models assume access to an empirical risk minimizer. In this work, we present a principled theoretical framework analyzing diffusion models, providing a state-of-the-art sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$. Our structured decomposition of the score estimation error into statistical and optimization components offers critical insights into how diffusion models can be trained efficiently. This analysis addresses a fundamental gap in the literature and establishes the theoretical tractability and practical relevance of diffusion models.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving impressive performance across tasks such as image synthesis, molecular design, and audio generation. Central to the training of these models is the estimation of the score function, which characterizes the reverse-time dynamics in the diffusion process. Diffusion models are widely adopted in computer vision and audio generation tasks (Ulhaq & Akhtar, 2022; Bansal et al., 2023), text generation (Li et al., 2022), sequential data modeling (Tashiro et al., 2021), reinforcement learning and control (Zhu et al., 2023), and life sciences (Jing et al., 2022; Malusare & Aggarwal, 2024). For a more comprehensive exposition of applications, we refer readers to survey paper (Chen et al., 2024).
While diffusion models exhibit strong empirical performance, understanding their sample complexity is essential to guarantee their efficiency, generalization, and scalability, enabling high-quality generation with minimal data in real-world, resource-constrained scenarios. Some of the key works studying the sample complexity are summarized in Table 1. A key limitation of sample complexity analyses of diffusion models done thus far is the lack of the presence of finite-time sample complexity results under reasonable assumptions. This makes the theoretical analysis of diffusion models fall short of other machine learning areas such as reinforcement Learning (Kumar et al., 2023; Gaur et al., 2024), bi-level-optimization (Grazzi et al., 2023; Gaur et al., 2025) and graphical models (Fattahi et al., 2019; Tran et al., 2019). In this work we aim to bridge that gap and obtain a sample complexity results on the same footing as results from the aforementioned areas. The iteration complexity or convergence has been studied in Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b), while they assume bounded score estimates thus not providing the sample complexity which requires estimating the score function.
@@
We note that works such as Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023); Chen et al. (2023) have sample complexity results that depend exponentially on the data dimension, making the result less useful in high-dimensional settings. Recently, Gupta et al. (2024) improved upon this by obtaining $\tilde{\mathcal{O}}(\epsilon^{-5})$ sample complexity without exponential dependence on data dimension. [^1] We note that Gupta et al. (2024) claimed a sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-3})$, while this claim does not account for the accumulation of errors across discretization steps. Specifically, their bound at each step depends on $\tilde{\mathcal{O}}(1/\epsilon^{2})$ samples, and applying a union bound over $\tilde{\mathcal{O}}(1/\epsilon^{2})$ steps yields a total sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-5})$. For more details, see Appendix A.. However, this work assumes access to the empirical risk minimizer (ERM) of the score estimation loss, a significant restriction that was explicitly highlighted as an open problem in Gupta et al. (2024) itself. While this assumption is present in all prior works, it is an unrealistic assumption regardless.
In this paper, we do not make this assumptions and establish an improved state-of-the-art sample complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-4})$. This represents a key step toward bridging the gap between the theory and practice of diffusion models. Specifically, we address the following fundamental question.
@@
How many samples are required for a sufficiently expressive neural network to estimate the score function well enough to generate high-quality samples using a DDPM algorithm?
Our analysis directly connects the quality of the learned score function to the total variation distance between the generated and target distributions, offering more interpretable and practically relevant guarantees. Additionally, our work accounts for the unavailability of the empirical risk minimizer. Using our novel analysis of the score estimation error, we obtain the sample complexity bounds without exponential dependence on the data-dimension. Our principled analysis accounts for the statistical and optimization errors while not assuming access to the empirical risk minimizer of the score estimation loss, and achieves state-of-the-art sample complexity bounds.
@@
Zhang et al. (2024) $\tilde{O}\left(\epsilon^{-d}\right)$ Yes
Wibisono et al. (2024) $\tilde{O}\left(\epsilon^{-(d)}\right)$ Yes
Oko et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Chen et al. (2023) $\tilde{O}\left(\epsilon^{-O(d)}\right)$ Yes
Gupta et al. (2024)1 $\tilde{O}\left(\epsilon^{-5}\right)$ Yes
This work $\tilde{O}(\epsilon^{-4})$ No
Table 1: Summary of sample complexity results for diffusion models, assuming no upper bound on score estimation error. For further details on how the sample complexity bounds are derived for Gupta et al. (2024), see Appendix A.
The statistical error occurs due to the finite sample size used to obtain the score estimate. Existing methods used to bound statistical errors assume bounded loss functions, which is not true in the case of diffusion models. We thus use a novel analysis that uses the conditional normality of the score function to obtain upper bounds for the statistical error.
Finally, the optimization error occurs due to a finite number of SGD steps during the estimation of the score function. It is precisely the error that was not accounted for in the previous works due to the assumption that they have access to the empirical risk minimizer. We use the quadratic growth property implied by the Polyak-Łojasiewicz (PL) condition assumed in Assumption 2 and a novel recursive analysis of the error at each stochastic gradient descent (SGD) step to upper bound this error.
@@
  • Finite time sample complexity bounds. We derive state-of-the-art sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-4})$ for score-based diffusion models, without exponential dependence on the data dimension or neural network parameters. Our analysis avoids the unrealistic assumptions used in prior works such as access to an empirical loss minimizer.
@@
Unconditional and Conditional Diffusion Models. Diffusion models have emerged as leading frameworks across vision, audio, and scientific domains. Foundational works such as Sohl-Dickstein et al. (2015) and Ho et al. (2020) introduced and refined Denoising Diffusion Probabilistic Models (DDPMs), enabling high-quality sample generation. Subsequent advances include improved noise schedules (Nichol & Dhariwal, 2021), score-based SDE formulations (Song et al., 2021), and efficient latent-space generation via Latent Diffusion Models (LDMs) (Rombach et al., 2022). Conditional diffusion models extend these techniques for guided generation tasks, with applications in time-series Tashiro et al. (2021), speech Huang et al. (2022), and medical imaging Dorjsembe et al. (2023). Conditioning mechanisms range from classifier-based Dhariwal & Nichol (2021) to classifier-free guidance Ho & Salimans (2022), which enabled text-to-image models like Imagen Saharia et al. (2022) and Stable Diffusion Rombach et al. (2022). Recent innovations focus on adaptive control Castillo et al. (2025), compositionality Liu et al. (2023), and multi-modal conditioning Avrahami et al. (2022).
Related Works: Despite the empirical success of diffusion models, theoretical understanding regarding the sample complexity remains limited. Assuming accurate score estimates, authors in Chen et al. (2022) showed that score based generative models can efficiently sample from a sub-Gaussian data distribution. Assuming a bounded score function, iteration complexity bounds have been extensively studied in recent works Li et al. (2024b); Benton et al. (2024); Li & Yan (2024); Huang et al. (2024); Dou et al. (2024); Liang et al. (2025a, b). In particular, Benton et al. (2024); Li et al. (2024b) establishes iteration complexity guarantees for DDPM algorithms. Several studies propose accelerated denoising schedules to improve these rates Li et al. (2024a); Liang et al. (2025b); Dou et al. (2024). Additionally, improved convergence rates under low-dimensional data assumptions are demonstrated in Li & Yan (2024); Huang et al. (2024); Liang et al. (2025a). In contrast to these works, our analysis addresses the sample complexity of score-based generative models, where the errors introduced by the neural network approximation, data sampling, and optimization are also accounted for.
Sample complexity bounds for diffusion models have been studied via diffusion SDEs under smoothness and spectral assumptions in Chen et al. (2023); Zhang et al. (2024); Wibisono et al. (2024); Oko et al. (2023). However, these bounds are exponential in the data dimension. Further, authors of Gupta et al. (2024) use the quantile-based approach to get the sample complexity bounds. In this work, we further improve on these guarantees. The detailed comparison of sample complexities with the key approaches mentioned above is provided in Table 1.
@@ 2 Preliminaries and Problem Formulation
A canonical forward process used in diffusion models is the Ornstein–Uhlenbeck (OU) process (Øksendal, 2003), defined by the following SDE
@@
where $B_{t}$ denotes standard Brownian motion. The solution of this SDE in closed form is given by
$x_{t}\sim e^{-t}x_{0}+\sqrt{1-e^{-2t}}{\epsilon},\quad\text{with }\epsilon\sim\mathcal{N}(0,I).$ (2)
@@
This is typically achieved using stochastic time-reversal theory (Anderson, 1982), which yields a corresponding reverse-time SDE as follows.
@@
where $\nabla\log p_{t}(x)$ is known as the score function of the distribution $p_{t}$. Simulating this reverse process starting from $x_{T}\sim p_{T}\approx\mathcal{N}(0,I)$ yields approximate samples from the original distribution $p_{0}$. This motivates a sampling strategy where we begin from $x_{T}\sim\mathcal{N}(0,I)$ for sufficiently large $T$, and then integrate the reverse SDE backward to $t=0$ using estimated score functions. In practice, the backward process is run up to a fixed time point $t_{0}$ known as the early stopping time and not $t=0$. This is done in order to improve performance and training speed (Lyu et al., 2022; Favero et al., 2025).
The continuous-time reverse SDE (Equation 3) is discretized over a finite sequence of times $0<t_{0}<t_{1}<\cdots<t_{K}=T$. The score function $s_{t}(x):=\nabla\log p_{t}(x)$ is approximated at these discrete points using a learned estimator $\hat{s}_{t_{k}}$. This discretization underlies the DDPM framework (Ho et al., 2020), where the reverse process is implemented by iteratively denoising the sample using the estimated scores at each time step. The detailed procedure is provided in Algorithm 1 in the Appendix B. We employ stochastic gradient descent (SGD) to learn the score function at each $t_{k}$, using either a constant or decreasing learning rate, as justified in our analysis later.
Problem formulation: Let the score function be approximated using a parameterized family of neural networks $\mathcal{F}_{\Theta}=\{s_{\theta}:\theta\in\Theta\}$, where each $s_{\theta}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ is represented by a neural network of depth $D$ and width $W$ with smooth activation functions. Given $n_{k}$ i.i.d. samples $\{x_{i}\}_{i=1}^{n}$ from the data distribution $p_{t_{k}}$, the score network is trained by minimizing the following time-indexed loss:
$\mathcal{L}_{k}(\theta):=\mathbb{E}_{x\sim p_{t_{k}}}\left[\|s_{\theta}(x,t_{k})-\nabla\log p_{t_{k}}(x)\|^{2}\right].$ (4)
Objective. Our goal is to quantify how well the learned generative model $\hat{p}_{t_{0}}$ approximates the true data distribution $p$ in terms of total variation (TV) distance. Specifically, we aim to show the number of samples needed so that with high probability, the TV distance $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$ is bounded by $\mathcal{O}(\epsilon)$, where $\epsilon$ is the $L^{2}$ estimation error of the score function. This reduces the generative performance analysis to establishing tight sample complexity bounds on the score estimation error, using only $n$ i.i.d. samples from $p$. We additionally define the following probability distributions:
$p_{t_{0}}:=$ Distribution obtained after backward process till time $t_{0}$ steps starting form $p_{T}$
$p^{dis}_{t_{0}}:=$ Distribution obtained by backward process till time $t_{0}$ starting from $p_{T}$
@@ 3 Sample Complexity of Diffusion Models
We first state the assumptions required throughout this work.
Assumption 1 (Bounded Second Moment Data Distribution.).
[…]
Some works that analyze the convergence of score-based diffusion models, such as Chen et al. (2022); Oko et al. (2023), assume that the data distribution is supported on a bounded set, thereby excluding commonly encountered distributions such as Gaussian and sub-Gaussian families. In contrast, our analysis only requires the data distribution to be sub-Gaussian, making our results applicable to a significantly broader class of distributions.
Assumption 2 (Polyak - Łojasiewicz (PL) condition.).
The loss $\mathcal{L}_{k}(\theta)$ for all $k\in[0,K]$ satisfies the Polyak–Łojasiewicz condition, i.e., there exists a constant $\mu_{t}>0$ such that
$\frac{1}{2}\|\nabla\mathcal{L}_{k}(\theta)\|^{2}\geq\mu_{t}\left(\mathcal{L}_{k}(\theta)-\mathcal{L}_{k}(\theta^{*})\right),\quad\forall~\theta\in\Theta,$ (5)
where $\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}_{k}(\theta)$ denotes the global minimizer of the population loss.
The Polyak-Łojasiewicz (PL) condition is significantly weaker than strong convexity and is known to hold in many non-convex settings, including overparameterized neural networks trained with mean squared error losses (Liu et al., 2022). Prior works such as Gupta et al. (2024) and Block et al. (2020) implicitly assume access to an exact empirical risk minimizer (ERM) for score function estimation, as reflected in their sample complexity analyses (see Assumption A2 in Gupta et al. (2024) and the definition of $\hat{f}$ in Theorem 13 of Block et al. (2020)). This assumption, however, introduces a major limitation for practical implementations, where exact ERM is not attainable.
Assumption 3 (Approximation error of the Class of Neural Networks).
For all $t\in[0,T]$, there exists a neural network parameter $\theta\in\Theta$ such that
[…]
This error is independent of the sampling algorithm, and describes the error due to neural network parametrization. In learning theory, it is common to treat the approximation error of a model class as a constant so that analyzes can focus on the estimation/ optimization terms dependent on the sample. This convention appears in standard excess-risk decompositions for fixed hypothesis classes (Shalev-Shwartz & Ben-David, 2014). In PAC-Bayesian analyses, approximation errors are denoted by a constant once the class is fixed (Mai, 2025). In (NTK/RKHS) analyses of neural networks, where it is assumed the target function lies in, or is well approximated by the specified function class, the misspecification error is represented as a constant term (Bing et al., 2025). In reinforcement learning algorithm analysis such as policy gradient, a task-dependent “inherent Bellman” or function-approximation error that remains constant while deriving performance rates (Mondal & Aggarwal, 2024; Fu et al., 2021; Gaur et al., 2024; Ganesh et al., 2025). Note that Gupta et al. (2024) also makes the same assumption implicitly, but assumes this constant to be zero (in Assumption A2).
Note that in certain works such as (Jiao et al., 2023), it is shown that the network size has to exponential in data dimension in order to achieve a small approximation error. However, in practice that would require an impractically large neural network size. In practice neural network size is of the same order as the data dimension. Thus for a fixed neural network size that we assume in this work, it makes sense to assume the approximation error as a constant.
@@
For all $k\in[0,K]$, the population loss $\mathcal{L}_{k}(\theta)$ is $\kappa$-smooth with respect to the parameters $\theta$, i.e., for all $\theta,\theta^{\prime}\in\Theta$
$\|\nabla\mathcal{L}_{k}(\theta)-\nabla\mathcal{L}_{k}(\theta^{\prime})\|\leq\kappa\|\theta-\theta^{\prime}\|.$ (7)
We assume that the estimators of the gradients ${\nabla}\mathcal{L}_{k}(\theta)$ have bounded variance.
@@
Together, these assumptions form a minimal yet sufficient foundation for analyzing score estimation in practice. Smoothness and bounded gradient variance implied by the sub-Gaussian assumption are mild and generally satisfied for standard neural architectures with ReLU or GELU activations and well-behaved data distributions. The PL condition has been shown to emerge in over-parameterized networks or under lazy training regimes, where the function class is expressive enough to approximate the ground-truth score function (Liu et al., 2022). Notably, these conditions are not only specific to our setting they have been widely adopted in recent works studying the optimization landscape of deep diffusion models (Salimans & Ho, 2022; Liu et al., 2022). Note that in no prior works were such assumptions stated since they assumed access to the empirical risk minimizer.
Theorem 1 (Total Variation Distance Bound).
Let $p_{t_{0}}$ denote the distribution obtained by the backward process till time $t_{0}$ starting form $p_{T}$, and $\hat{p}_{t_{k}}(x)$ be the distribution generated by the backward process at discretized time steps $\{t_{k}\}$ using the estimated score functions $\hat{s}_{T-t_{k}}(x)$ where $k\in[0,K]$. Let $d$ be the data dimension, and $n_{k}$ be the number of samples for score estimation at time step $t_{k}$.
Assume that the data distribution satisfies Assumption 1, the loss function $\mathcal{L}_{k}(\theta)$ satisfies Assumptions 2 3,4 for all $k\in[0,K]$ and the learning rate for the $i^{th}$ SDG step with one sample per step, satisfies $\eta_{i}=\frac{\alpha}{i+\gamma}$ where $\alpha.\mu>1$ and $\gamma>\alpha.\kappa$. Further assume
$n_{k}={\Omega}\left(W^{2D}{\cdot}d^{2}.\log\left(\frac{4K}{\delta}\right)\left(\frac{\epsilon^{-4}}{\sigma_{k}^{-4}}\right)\right),$ (9)
Then, with probability at least $1-\delta$, the total variation distance between the $p_{t_{0}}$ and $\hat{p}_{t_{0}}$ satisfies
$TV(p_{t_{0}},\hat{p}_{t_{0}})\leq\mathcal{O}(\exp^{-T})+\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)+\mathcal{O}(\epsilon)+\epsilon_{approx}$ (10)
Furthermore, by setting $T=\Omega\left(\log\left(\frac{1}{\epsilon}\right)\right)$ and $K=\Omega(\epsilon^{-2})$, we obtain
$\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})\leq{\mathcal{O}}(\epsilon)+\epsilon_{approx},$ (11)
with probability at least $1-\delta$.
Theorem 1 establishes that the total variation distance between the true data distribution and the diffusion model’s output can be made arbitrarily small specifically, $\tilde{\mathcal{O}}(\epsilon)$ by properly scaling model capacity and algorithmic parameters. To the best of our knowledge, these are the only known sample complexity bounds for score-based diffusion models, improving upon the prior results as discussed in the introduction without assuming access to empirical risk minimizer for the score estimation loss. Our results highlight an inherent trade-off between the sample complexity bounds and the assumption of access to the empirical risk minimizer.
Usage of $p_{t_{0}}$ instead of $p_{0}$ in Theorem 1: We have shown that the estimated distribution $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in total variation (TV) to $p_{t_{0}}$, where $p_{t_{0}}$ denotes the data distribution pushed forward by $t_{0}$ steps of the forward process. We do not claim that $\hat{p}_{t_{0}}$ is $\mathcal{O}(\epsilon)$-close in TV to the true data distribution $p_{0}$ (i.e., we do not bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$), because doing so would require additional assumptions on $p$. For example, Fu et al. (2024) (in Lemma D.5) assumes a sub-Gaussian data distribution to show that $\mathrm{TV}(p_{0},p_{t_{0}})\leq\mathcal{O}\left(\sqrt{t_{0}}\log\!\bigl(1/t_{0}\bigr)\right)$. We also note that all other works listed in Table 1 similarly provide upper bounds on $\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}})$, not on $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$.
However, it is to be noted that using the sub-Gaussian assumption, our analyis can be extended to a bound $\mathrm{TV}(p_{0},\hat{p}_{t_{0}})$ via the triangle inequality:
$\mathrm{TV}(p_{0},\hat{p}_{t_{0}})\;\leq\;\mathrm{TV}(p_{0},p_{t_{0}})\;+\;\mathrm{TV}(p_{t_{0}},\hat{p}_{t_{0}}).$
We formally present the data assumption and the resulting theorem as follows